OpenStudy (alexa):

What is the missing step in this proof? 1.∠ABC ≅ ∠DBE 2.∠BCA ≅ ∠BDE 3.∠ACB ≅ ∠DEB 4.∠BDE ≅ ∠ADE 5.∠CAB ≅ ∠DAC

OpenStudy (alexa):

1.AD/DB=CE/EB GIVEN 2.AD/DB+=CE/EB+1 ADDITION PROPERTY OF EQUALITY 3.AD+DB/DB=CE+EB/EB USING COMMON DENOMINATORS 4.AB=AD+DB and CB=CE+EB SEGMENT ADDITION 5.AD/DB=CE/DB SUBSTITUTION PROPERTY OF EQUALITY 6. REFLEXIVE PROPERTY OF CONGRUENCE 7.triangle ABC ~ triangle DBE SAS SIMILARITY CRITERION 8.angle BAC=~angle BDE CORRESPONDING ANGLES OF SIMILAR TRIANGLES ARE CONGRUENT 9.DE//AC If the corresponding angles formed by two lines cut by a transversal are congruent, then the lines are parallel.

OpenStudy (alexa):

is it the last one?

OpenStudy (phi):

I would expect its the one that shows the "shared angle" <DBE equals itself (but called <ABC )

OpenStudy (alexa):

oh the 4th one? Yea i might have had a typo but i cant go back to the question its online, if i go over it again, ill tell you what up, thanks for taking the time to answer!

OpenStudy (phi):

it would be choice 1

OpenStudy (alexa):

OHHH what, oh wow i never thought about that one but thank you @phi